# Alleviating Non-identifiability: a High-fidelity Calibration Objective for Financial Market Simulation with Multivariate Time Series Data

### v2

@@ Abstract
The non-identifiability issue has been frequently reported in the social simulation works, where different parameters of an agent-based simulation model yield indistinguishable simulated time series data under certain discrepancy metrics. This issue largely undermines the simulation fidelity yet lacks dedicated investigations. This paper theoretically analyzes that incorporating multiple time series data features in the model calibration phase can alleviate the non-identifiability exponentially with the increasing number of features. To implement this theoretical finding, a maximization-based aggregation function is applied to existing discrepancy metrics to form a new calibration objective function. For verification, the financial market simulation, a typical and complex social simulation task, is considered. Empirical studies on both synthetic and real market data witness the significant improvements in alleviating the non-identifiability with much higher simulation fidelity of the chosen agent-based simulation model. Importantly, as a model-agnostic method, it achieves the first successful simulation of the high-frequency market at seconds level. Hence, this work is expected to provide not only a rigorous understanding of non-identifiability in social simulation, but a high-fidelity calibration objective function for financial market simulations.
@@ 1 Introduction
Social simulation plays a pivotal role of understanding the complex social systems in a way of revealing their endogenous complexity  [1]. The core is to construct a parameterized simulation model that can generate the simulated time series data resembling the observed states over time of the social system. By properly modeling the endogenous components of the social system as multiple interacting agents, i.e., the Agent-based Modeling (ABM) techniques  [2], various “what-if” tests can be conducted  [3]. That is, by intervening certain agents of the simulation model, the corresponding endogeneity of the real social system is revealed by observing how the evolution of the simulation model is influenced by the intervention.
@@
The success of a social simulation largely lies with the model fidelity. Note that, the what-if tests are often counterfactual, i.e., they do not happen in real social systems and thus there are no ground-truth for validation. To what extent should we believe the results of what-if tests? In social systems, due to the uncertainty of human activities, it is often impractical to derive trustworthy simulation models mathematically or physically. Usually, this issue is addressed in a data-driven manner via model calibration  [4].
@@
The basic idea is that, if a simulation model $M(\omega)$ can generate simulated data $\mathbf{X}_{T}$ resembling the observed data $\hat{\mathbf{X}}_{T}$ of the targeted real system within the discrete time interval $T=[t_{s},t_{s}+1,\cdots,t_{e}]$, it is considered that the underlying probabilistic distribution of the real system for generating $\hat{\mathbf{X}}_{T}$ is learned by $M(\omega)$. Here, $\hat{\mathbf{X}}_{T}$ denotes the time series $\{\hat{\mathbf{X}}_{t}\}^{t=t_{e}}_{t_{s}}$ for short. Hence, while using $M(\omega)$ for what-if test, by keeping the non-intervened agents of $M(\omega)$ unchanged, the simulation results of how the intervened agent influences the system are considered trustworthy. To obtain such $M(\omega)$, the discrepancy between $\mathbf{X}_{T}$ and $\hat{\mathbf{X}}_{T}$ is measured by various metrics $D$  [5]. The smaller $D$ is, the closer $M(\omega)$ approximates the distribution of data generation behind the real social system. By fixing the ABM-based model structure $M$, calibrating the parameter $\omega$ to fit a given observed data sequence $\hat{\mathbf{X}}_{T}$ is often modeled as an optimization problem  [5]:
$\min_{\omega\in\Omega}D(\hat{\mathbf{X}}_{T},M(\omega)=\mathbf{X}_{T})$ (1)
@@
In practice, the above data-driven calibration methods often suffer from the so-called non-identifiability issue  [6, 7]. That is, there are multiple parameters in the parameter space $\Omega$ enjoy the same discrepancy value, given the same $\hat{\mathbf{X}}_{T}$. However, those parameters highly likely lead to different distributions of data generation, due to the nonlinear model structure $M(\omega)$. In other words, the “ground-truth” parameter, deciding the “ground-truth” data generation distribution, cannot be effectively identified from the other parameters in those sets. Consequently, the what-if tests on the resultant model of a randomly picked non-identifiable parameter therefrom are less trustworthy.
@@
Despite that the non-identifiability is frequently mentioned in the literature, there is neither a formal definition of it nor a rigorous way to alleviate it. This paper first defines the non-identifiability issue in the form of multivariate time-series data. It is then discussed that the non-identifiability arises from that the previously calibrated time series data $\hat{\mathbf{X}}_{T}$ is often univariate and is insufficient to describe the uniqueness of a social system. For simplicity and applicability, this work restricts the scenarios to the financial market simulation, as it is a representative social simulation and has been active for over 30 years  [8]. Like other complex social systems, the real-world financial markets will output multivariate states data over time, e.g., price, volume, bid/ask directions, order arrival time, and other hand-crafted features  [9]. However, existing works merely consider the observed price data for calibration, resulting in wide criticism for their low simulation fidelity  [10], especially in intraday simulations  [11, 6].
This paper alleviates the non-identifiability issue by incorporating more data features into the calibration process, each of which represents a univariate time series data. We mathematically show that the non-identifiability can be reduced exponentially with the number of features. To this end, the multivariate time series is fed into existing discrepancy metrics in a way of multiple univariate time series. That is, the data associated with each feature is calibrated individually. This results in multiple individual calibration tasks with respect to the multiple features, respectively. Note that, the non-identifiable set of the multivariate time series calibration task should be non-identifiable across all the individual tasks. Specifically, only the ground-truth parameter locates simultaneously at the global optimum of all parameter spaces of those individual tasks. To implement it, we propose to search in the intersection of all those individual spaces and implement it via a maximization function aggregating all those individual calibration tasks. Such an implementation enjoys a high scalability that one can incorporate as many as possible data features to alleviate the non-identifiability.
@@
Extensive empirical studies have been conducted to show that the proposed method can significantly alleviate the non-identifiability issue of financial market simulation on 10 synthetic data and 1 real data from Shenzhen Stock Exchange of China. Such advantages are verified robust over 6 commonly seen financial market data features. Besides, this paper successfully reports the high-fidelity simulation results of the high-frequency market at seconds level, which is the first time for the ABM-agnostic calibration methods.
@@
The rest of this paper is organized as follows. Section  2 describes the background of financial market simulation and related works. Section  3 theoretically discuss the alleviation of non-identifiability. Section  4 reports the empirical studies in details. Section  5 concludes this work.
@@ 2 Background
2.1 ABM-based Financial Market Simulation
In social simulations, Agent-based Modeling (ABM) based models are dominantly preferred than the black-box deep learning models due to their interpretability of how the data is generated  [12]. ABM constructs a collection of agents mimicking various social activities and interacting together to simulate the whole social system  [12]. In financial market simulation, the agents are designed to model various trading strategies of human  [13] (illustrated in Figure  1), e.g., fundamentalists, momentum traders, high-frequency traders, market makers, and arbitrage traders. Recent financial market simulation works favor the Preis-Golke-Paul-Schneid (PGPS) model, which consists of two general types of agents, i.e., the liquidity providers and the liquidity takers  [14].
Figure 1: An illustration of the ABM-based financial market simulator with various agents interacting with the simulated exchange.
Regardless of the differences among agents, the ABM-based financial market simulations follow the same workflow. That is, at the beginning of the simulation, the parameterized agents are initialized with some parameters, which determine how exactly those agents submit orders under their own simulated strategies. Each submitted order usually consists of the price, volume, direction, and arrival time. The exchange is simulated as a Limited Order Book (LOB) that maintains the existing untraded orders submitted by the agents  [15].
Figure 2: A LOB organizes all the untraded orders in terms of directions, prices, arrival time and volume.
Those untraded orders are categorized into two directions (bid or ask) and sorted by their prices. For the orders with the same price, they are again sorted by the arrival time. Each agent continuously observes the current LOB and submit its order at will. If a new order matches a price in the opposite direction of the LOB, it will be traded immediately, and the matched orders will be removed from the LOB  [16]. In existing financial market simulations, the mid-price (see Figure  2) is mostly considered as the observable time series data for calibration, as it implies the potential movement of the trading prices and is important to many what-if tests, e.g., the causes of flash crashes  [17, 18].
2.2 Calibration of Financial Market Simulation
@@
To ensure the fidelity of a financial market simulation, the model often needs careful calibration of the agents’ parameters, so that the simulated data resembles the observed time series from the real market. The general form of calibration has been given in Eq. (1).
@@
In recent years, the study of the discrepancy metrics $D$ has attracted increasing research attentions, as the distance measure between time series remains unsolved and impacts significantly on the calibration performance. Methods of Simulated Moments  [19] are a typical method that first transforms the time series into several statistical moments, and then calculates the weighted average between the moments of observed data and simulated data. The information criterion based objective functions are also studied  [20, 21]. These methods abstract the temporal information of the original time series via various techniques like histograms and measure the distances between the abstracted temporal vectors. The non-parametric Kolmogorov-Smirnov test is also introduced as $D$ by measuring the distance between the probabilistic distributions of the observed data and the simulated data  [7]. Others employ the tool of Bayesian theory to estimate the likelihood  [22] or posterior distribution  [23] of the observed data.
@@
Despite of the progresses of exploiting various metrics for calibrating the financial market simulation, existing works frequently encounter the so-called non-identifiability issue, resulting in poor simulation fidelity  [24]. From our viewpoint, one major reason is that existing works merely use the mid-price data to calibrate the model, ignoring that the real financial market continuously output the data with multiple features. Although the mid-price might be important to the market participants, the other features describe the market in more details. Unfortunately, the above mentioned calibration objective functions are all designed for univariate time series data (for mid-price). Consequently, existing works are ill-equipped to calibrate with multivariate time series data. Yet, it is still unclear how exactly it benefits from calibrating more features.
@@ 3 A High-fidelity Calibration Objective
This section first formally defines the non-identifiability, then theoretically analyzes how it benefits from calibrating with multivariate data, and lastly proposes a new calibration objective function using multiple features.
@@ 3.1 The Definition of Non-identifiability
As existing discrepancy metrics $D$ are designed for univariate data, we re-denote the calibrated data as $\hat{\mathbf{X}}_{T}^{k}$ with a superscript $k$ to emphasize that it is a 1-dimensional time series describing some $k$-th feature of the real system. In financial markets, this feature can be either price, volume, bid/ask directions, order arrival time, or any other hand-crafted features  [9]. Correspondingly, the simulated data generated by $M(\omega)$ is re-denoted as $\mathbf{X}^{k}_{T}$. Then the original calibration task can be re-written as $\min_{\omega}D(\hat{\mathbf{X}}_{T}^{k},M(\omega)=\mathbf{X}^{k}_{T})$, and the non-identifiability issue is defined as follows.
Let $\Omega^{D,k}$ be the parameter space of this calibration task. There exists some region $\mathcal{S}^{D,k}\subseteq\Omega^{D,k}$ that any two different parameters $\omega_{1},\omega_{2}\in\mathcal{S}^{D,k}$ satisfy $D(\hat{\mathbf{X}}_{T}^{k},M(\omega_{1}))=D(\hat{\mathbf{X}}_{T}^{k},M(\omega_{2}))$. Such region $\mathcal{S}^{D,k}$ is called the non-identifiability set and is non-empty if the ground-truth parameter $\omega^{*}$ is assumed to be included in $\Omega^{D,k}$. Then the non-identifiability is defined as the probability of uniformly randomly sampling a parameter $\omega$ from $\Omega^{D,k}$ while falling into $\mathcal{S}^{D,k}$, denoted as $P(\omega\in\mathcal{S}^{D,k}|\hat{\mathbf{X}}_{T}^{k})$. Let $|\mathcal{S}|$ denote the cardinality of a set $\mathcal{S}$, it gives
$P(\omega\in\mathcal{S}^{D,k}|\hat{\mathbf{X}}_{T}^{k})=\frac{|\mathcal{S}^{D,k}|}{|\Omega^{D,k}|}$ (2)
Therefore, the rank of any $\omega_{1},\omega_{2}\in\mathcal{S}^{D,k}$ cannot be effectively identified by the objective $D$ given the data $\hat{\mathbf{X}}_{T}^{k}$, leading to a flat landscape of the parameter space therein and the search-based calibration process easily gets stuck in $\mathcal{S}^{D,k}$. Note that, $\mathcal{S}^{D,k}$ can involve arbitrary parameters in $\Omega^{D,k}$, including the “ground-truth” parameter $\omega^{*}$. Hence, if the calibration process converges to $\mathcal{S}^{D,k}$, one has to randomly pick one $\omega\in\mathcal{S}^{D,k}$ for further what-if simulation. Unfortunately, the underlying data generation distribution decided by any $\omega\in\mathcal{S}^{D,k}$ is less likely to be the same with $\omega^{*}$, given the nonlinear nature of the simulation model $M(\omega)$. Consequently, the what-if test on the randomly picked $\omega$ is less trustworthy.

### v3

@@ Abstract
The non-identifiability issue has been frequently reported in the social simulation works, where different parameters of an agent-based simulation model yield indistinguishable simulated time series data under certain discrepancy metrics. This issue largely undermines the simulation fidelity yet lacks dedicated investigations. This paper theoretically analyzes that incorporating multiple time series data features in the model calibration phase can alleviate the non-identifiability exponentially with the increasing number of features. To implement this theoretical finding, a maximization-based aggregation function is proposed based on existing discrepancy metrics to form a new calibration objective function. For verification, the task of calibrating the Financial Market Simulation (FMS), a typical yet complex social simulation, is considered. Empirical studies confirm the significant improvements in alleviating the non-identifiability of calibration tasks. Furthermore, as a model-agnostic method, it achieves much higher simulation fidelity of the chosen FMS model on both synthetic and real market data. Hence, this work is expected to provide not only a rigorous understanding of non-identifiability in social simulation, but an off-the-shelf high-fidelity calibration objective function for FMS.
@@ 1 Introduction
Social simulation plays a pivotal role of understanding the complex social systems in a way of revealing their endogenous complexity  [27]. By properly modeling the endogenous components of the social system as multiple interacting agents, i.e., using the Agent-based Models (ABMs)  [10], various “what-if” tests can be conducted by intervening the agents of interest and analyzing the evolution process of the simulators  [8, 29].
@@
The success of a social simulation largely lies with the model fidelity. Note that, the what-if tests are often counterfactual, i.e., they do not happen in real social systems and thus there are no ground-truth for validation. To what extent should we believe the results of what-if tests? Besides, due to the uncertainty of human activities, it is often impractical to derive mathematically trustworthy social simulation models. Usually, this issue is addressed in a data-driven manner via model calibration  [16].
@@
Specifically, a simulation model $M(\omega)$ is essentially a data generative process, i.e., $M(\omega)=\mathbf{X}_{T^{\prime}},T^{\prime}=[1,...,\tau],\tau\in\mathbb{N}^{+}$. If the simulated data $\mathbf{X}_{T}$ resembles the observed data $\hat{\mathbf{X}}_{T}$ of the targeted real system within any discrete time interval of interest $T=[t_{s},t_{s}+1,...,t_{e}]$, it is considered that the underlying data generating probabilistic distribution of the real system for generating $\hat{\mathbf{X}}_{T}=\{\hat{\mathbf{X}}_{t}\}^{t_{e}}_{t=t_{s}}$ is learned by $M(\omega)$. Hence, while using $M(\omega)$ for what-if test, by keeping the non-intervened agents of $M(\omega)$ unchanged, the simulation results of how the intervened agent influences the system are considered trustworthy. To obtain such $M(\omega)$, the discrepancy between $\mathbf{X}_{T}$ and $\hat{\mathbf{X}}_{T}$ is measured by various metrics $D$  [24]. The smaller $D$ is, the closer $M(\omega)$ approximates the data generating distribution of the real social system. By fixing the ABM-based model structure $M$, calibrating the parameter $\omega$ to fit a given observed data sequence $\hat{\mathbf{X}}_{T}$ is often modeled as an optimization problem  [24]:
$\min_{\omega\in\Omega^{D}}D(\hat{\mathbf{X}}_{T},M(\omega)=\mathbf{X}_{T})$ (1)
@@
In practice, the calibration methods often suffer from the so-called non-identifiability issue  [11, 1]. That is, there are multiple parameters in the parameter space $\Omega^{D}$ share the same discrepancy value to the given $\hat{\mathbf{X}}_{T}$. However, those parameters highly likely lead to different distributions of data generation, due to the nonlinear model structure $M(\omega)$. In other words, the “ground-truth” parameter of deciding the “ground-truth” data generating distribution cannot be effectively identified from the other parameters in those sets. Consequently, the what-if tests on the resultant model of a randomly picked non-identifiable parameter therefrom are less trustworthy.
Despite that the non-identifiability is frequently mentioned in the literature, there is neither a formal definition to describe it nor a rigorous way to alleviate it. This paper first defines the non-identifiability as the probability of a randomly uniformly sampled parameter falling into the non-identifiable set. With this definition, this paper mathematically proves that the non-identifiability can be alleviated by calibrating with more distinct features, each of which represents an observed univariate time series. More specifically, we mathematically show that the upper bound of the non-identifiability can be reduced exponentially with the number of features, if the selected features satisfy certain conditions.
@@
This explains well that the previous works mostly calibrate their models to merely one data feature and is thus less sufficient to describe the uniqueness of a social system. For clarification and applicability, this work restricts the scenarios to the Financial Market Simulation (FMS), as it is a representative social simulation and has been active for over 30 years  [22]. Like other complex social systems, the real-world financial markets will output multivariate states data over time, e.g., price, volume, bid/ask directions, order arrival time, and other hand-crafted features  [21]. However, existing works merely consider the observed mid-price data for calibration, resulting in wide criticism for their low simulation fidelity  [9], especially in the high-frequency intraday simulations  [25, 11].
Note that, following our theory, the task of calibrating to multivariate time series is naturally divided into multiple individual tasks, each of which calibrates the model to one distinct univariate time series. In practice, how to jointly execute those individual tasks so that the obtained parameter is more identifiable to the original calibration tasks with multivariate time series? This challenges the implementation of the above theory. To address it, we propose a novel calibration objective by aggregating all those individual calibration tasks using the maximization function. This new objective function mathematically asks to search in the intersection of all those individual spaces, thus the obtained parameter can be highly identifiable across all individual tasks and thus fit our theory accurately.
@@
Extensive empirical studies have successfully shown that the proposed method can significantly alleviate the non-identifiability issue of financial market simulation. Such advantages are verified robust over 6 commonly seen financial market data features. Besides, this paper also presents the high-fidelity simulation results by using the proposed new calibration objective function on 10 synthetic data and 1 real data from Shenzhen Stock Exchange of China.
@@
The reminder of this paper is as follows. Section 2 describes the background of financial market simulation and related works. Section 3 theoretically discuss the alleviation of non-identifiability. Section 4 reports the empirical studies in details. Section 5 concludes this work.
@@ 2 Background
2.1 Calibration of Social Simulators
@@
To ensure the simulation fidelity, the simulator often needs careful calibration of the agent parameters, so that the simulated data resembles the observed time series from the real system. The general form of calibration refers to Eq. (1).
@@
In recent years, the study of the discrepancy metrics $D$ has attracted increasing research attentions, as the distance measure between time series remains unsolved and impacts significantly on the calibration performance. Methods of Simulated Moments  [12] are a typical method that first transforms the time series into several statistical moments, and then calculates the weighted average between the moments of observed data and simulated data. The information criterion based objective functions  [18, 2] abstract the temporal information of the original time series via various techniques like histograms and measure the distances between the abstracted temporal vectors. The non-parametric Kolmogorov-Smirnov test is also introduced as $D$ by statistically testing between the probabilistic distributions of the observed data and the simulated data  [1]. Other methods employ the Bayesian theory to estimate the likelihood  [17] or posterior distribution  [13] of the observed data also require a well-defined $D$ for selecting effective samples and updating the estimated distributions.
Figure 1: An illustration of the ABM-based FMS.
@@
Despite of the progresses of exploiting various metrics for calibrating social simulators, existing works frequently encounter the so-called non-identifiability issue, resulting in poor simulation fidelity  [19]. From our viewpoint, one major reason is that existing works merely use the one single feature to calibrate the model, ignoring that the real systems continuously output multivariate data. Although the used feature might be a major factor to describe the social system, the other features can help capture more details of the system. Unfortunately, existing calibration objectives are mostly designed for univariate time series data. Consequently, existing works are ill-equipped to calibrate with multivariate time series data. Yet, it is theoretically unclear how exactly it benefits from calibrating more features.
2.2 ABM-based FMS and Data Features
In social simulations, ABMs are dominantly preferred than the black-box deep learning models due to their interpretability of the data generating process  [14, 20]. ABM constructs a collection of agents mimicking various social activities and interacting together to simulate the whole social system  [23]. In FMS, the agents are designed to model various trading strategies  [4] (illustrated in Figure 1), e.g., fundamentalists, momentum traders, high-frequency traders, market makers, and arbitrage traders.
Regardless of the differences among agents, the ABM-based FMSs follow the same workflow. That is, at the beginning of the simulation, the parameterized agents are initialized with the calibrated parameters, which determine how exactly those agents submit orders under their own simulated strategies. Each submitted order usually consists of the price, volume, direction, and arrival time. The Limited Order Book (LOB), a special data structure for organizing the existing untraded orders, is maintained by the exchange and can be observed by all trading agents. The simulated exchange generates the next LOB by matchmaking the incoming orders submitted by the agents and the current LOB  [28].
Those untraded orders are categorized into two directions (bid or ask) and sorted by their prices. For the orders with the same price, they are again sorted by the arrival time. Each agent continuously observes the current LOB and submit its order at will. If a new order matches a price in the opposite direction of the LOB, it will be traded immediately, and the matched orders will be removed from the LOB. Undoubtedly, many data features can be extracted from LOB  [21]. However, existing FMSs mostly considered the mid-price as the only feature for calibration, which represents the average between the best ask price and best bid prices (see Figure 2) and implies the potential price movement of the market.
Figure 2: LOB contains much information about the market.
@@ 3 A High-fidelity Calibration Objective
This section first formally defines the non-identifiability, then theoretically analyzes how it benefits from calibrating with multivariate data, and lastly proposes a new calibration objective using multiple features to implement our theory.
@@ 3.1 The Definition of Non-identifiability
As existing discrepancy metrics $D$ are designed for univariate data, we re-denote the target data as $\hat{\mathbf{X}}_{T}^{k}$ with a superscript $k$ to emphasize that it is a 1-dimensional time series describing some $k$-th feature of the real system. In financial markets, this feature can be either price, volume, bid/ask directions, order arrival time, or any other hand-crafted features  [21]. Correspondingly, the $k$-th feature of the simulated data is re-denoted as $\mathbf{X}^{k}_{T}$. Let $\Omega^{D,k}$ be the parameter space of this calibration task. Then the original calibration task can be re-written as
$\min_{\omega\in\Omega^{D,k}}D(\hat{\mathbf{X}}_{T}^{k},M(\omega)=\mathbf{X}^{k}_{T}).$ (2)
Without loss of generality, it is assumed that the optimal parameter $\omega^{*}$ lies in the defined parameter space, i.e., $\omega^{*}\in\Omega^{D,k}$, where $D(\hat{\mathbf{X}}_{T}^{k},M(\omega^{*}))<\epsilon$. Note that, the optimal parameter may not be unique by definition. The cardinality of the optimal set is influenced by $\epsilon$, which is used to filter the observation noise of data. Then the non-identifiable set is defined as follow:
Definition 1 (The non-identifiable set)
There exists some region $\mathcal{S}^{D,k,\epsilon}\subseteq\Omega^{D,k}$ that for all $\omega_{1},\omega_{2}\in\mathcal{S}^{D,k,\epsilon}$ such that $D(\hat{\mathbf{X}}_{T}^{k},M(\omega_{1}))<\epsilon$ and $D(\hat{\mathbf{X}}_{T}^{k},M(\omega_{2}))<\epsilon$. Such region $\mathcal{S}^{D,k,\epsilon}$ is called the non-identifiable set.
The cardinality of $\mathcal{S}^{D,k,\epsilon}$, denoted as $|\mathcal{S}^{D,k,\epsilon}|$, is also influenced by $\epsilon$. Then the non-identifiability can be defined as follows:
Definition 2 (The univariate non-identifiability)
Let $P(\omega\in\mathcal{S}^{D,k,\epsilon}|\hat{\mathbf{X}}_{T}^{k})$ be the non-identifiability to the $k$-th univariate time series data. It it defined as the probability of uniformly randomly sampling a parameter $\omega$ from $\Omega^{D,k}$ while it falls into $\mathcal{S}^{D,k,\epsilon}$. And we have
$P\left(\omega\in\mathcal{S}^{D,k,\epsilon}\middle|\hat{\mathbf{X}}_{T}^{k}\right)=\frac{|\mathcal{S}^{D,k,\epsilon}|}{|\Omega^{D,k}|}.$ (3)
Therefore, the rank of any $\omega_{1},\omega_{2}\in\mathcal{S}^{D,k,\epsilon}$ cannot be effectively identified by the objective $D$ given the data $\hat{\mathbf{X}}_{T}^{k}$, leading to that the comparison-based calibration process easily gets stuck in $\mathcal{S}^{D,k,\epsilon}$. Note that, $\mathcal{S}^{D,k,\epsilon}$ can involve arbitrary parameters in $\Omega^{D,k}$, including the “ground-truth” parameter $\omega^{*}$. Hence, if the calibration process converges to $\mathcal{S}^{D,k,\epsilon}$, one has to randomly pick one $\omega\in\mathcal{S}^{D,k,\epsilon}$ for further what-if simulation. Unfortunately, the underlying data generative distribution decided by any randomly picked $\omega\in\mathcal{S}^{D,k,\epsilon}$ is less likely to be the same with $\omega^{*}$, given the nonlinear nature of the simulation model $M(\omega)$. Consequently, the what-if test on the randomly picked $\omega$ is less trustworthy, unless the cardinality of $\mathcal{S}^{D,k,\epsilon}$ is sufficiently small.